OpenStudy (anonymous):

Fan and medal. A box of fewer than 100 but more than 12 cookies can be shared equally among 4, 10, or 12 people with no cookies left over. How many cookies are in the box?

OpenStudy (lucaz):

12 is divisible by 4, so anything divisible by 12 is also divisible by 4. you need a number divisible by 12 and 10. try 60 60/10 = 6 60/12 = 5 60/4 = 30

OpenStudy (anonymous):

Thats wrong..

OpenStudy (lucaz):

do you have the answer?